# InstructRAG: Instructing Retrieval-Augmented Generation via Self-Synthesized Rationales

### v2

@@ Abstract
Retrieval-augmented generation (RAG) has shown promising potential to enhance the accuracy and factuality of language models (LMs). However, imperfect retrievers or noisy corpora can introduce misleading or even erroneous information to the retrieved contents, posing a significant challenge to the generation quality. Existing RAG methods typically address this challenge by directly predicting final answers despite potentially noisy inputs, resulting in an implicit denoising process that is difficult to interpret and verify. On the other hand, the acquisition of explicit denoising supervision is often costly, involving significant human efforts. In this work, we propose InstructRAG, where LMs explicitly learn the denoising process through self-synthesized rationales — First, we instruct the LM to explain how the ground-truth answer is derived from retrieved documents. Then, these rationales can be used either as demonstrations for in-context learning of explicit denoising or as supervised fine-tuning data to train the model. Compared to standard RAG approaches, InstructRAG requires no additional supervision, allows for easier verification of the predicted answers, and effectively improves generation accuracy. Experiments show InstructRAG consistently outperforms existing RAG methods in both training-free and trainable scenarios, achieving a relative improvement of 8.3% over the best baseline method on average across five knowledge-intensive benchmarks. Extensive analysis indicates that InstructRAG scales well with increased numbers of retrieved documents and consistently exhibits robust denoising ability even in out-of-domain datasets, demonstrating strong generalizability.111Our code is available at https://github.com/weizhepei/InstructRAG.
@@ 1 Introduction
While large language models (LMs) have demonstrated remarkable text generation abilities [6, 81, 82], they may occasionally produce factually incorrect contents [15, 25, 31, 79, 99, 110], particularly when the task at hand requires the most current information or out-of-domain knowledge not adequately represented in the pre-training corpus [35, 75, 106, 113]. This limitation significantly hinders the reliable deployment of LMs in high-stakes domains where factuality is crucial [60, 76, 93, 94].
@@
In light of this, retrieval-augmented generation (RAG) [3, 23, 29, 42, 52] has been introduced to enhance the generation accuracy of LMs in knowledge-intensive tasks by leveraging the most up-to-date information and specialized knowledge from external sources [40, 83, 102, 117]. However, the retrieved contents are typically mixed with irrelevant or even erroneous information due to the absence of perfect retrieval solutions [27, 39, 43, 44, 74] and the presence of noisy data in the retrieval corpus [28, 53, 105], posing a long-standing challenge to almost all RAG systems. Typically, vanilla RAG approaches address this issue implicitly by training LMs to directly predict correct answers despite potentially noisy inputs. Such latent processes are not only difficult to interpret and verify but also vulnerable to higher noise ratios, especially when the number of retrieved documents is large [11, 13, 56, 91]. On the other hand, obtaining high-quality explicit denoising supervision often requires substantial human efforts, which is time-consuming and costly.
@@
In this work, we introduce a new RAG framework, InstructRAG, which enables the LM to explicitly denoise retrieved information and justify its predicted final answers by generating denoising responses (i.e., rationales), as illustrated in Figure 1. Compared to vanilla RAG approaches, InstructRAG does not require any additional supervision, while enjoying improved generation accuracy and trustworthiness. Specifically, our method consists of two steps. First, given a set of question-answer pairs and potentially noisy retrieved documents, we prompt an instruction-tuned LM to synthesize denoising rationales that analyze the documents and articulate how they lead to the ground-truth answers (§ 2.2). Then, these synthetic rationales can be utilized as in-context learning examples or as supervised fine-tuning data, allowing the LM to explicitly learn to denoise retrieved contents (§ 2.3). The effectiveness of InstructRAG can be attributed to the strong instruction-following ability of LMs [34, 67, 88], a significant feature that still remains underexplored in the context of RAG. We show that such self-synthesized rationales not only provide high-quality explicit denoising supervision for in-domain RAG tasks, but also facilitate superior out-of-domain generalization. This finding underscores how instruction-tuned LMs can synthesize generalizable supervision to effectively overcome the inevitable noise in RAG.
@@ 2 Our Method: InstructRAG
In this section, we first introduce our problem setting (§ 2.1) and then present the proposed framework InstructRAG that enables LMs to explicitly denoise retrieved contents. As shown in Figure 2, our method consists of two steps. First, we prompt an instruction-tuned LM (i.e., rationale generator ) to synthesize rationales that provide denoising supervisions (§ 2.2). These rationales aim to explain how to derive the correct answer from potentially noisy retrieved documents for each training sample. Then, we guide the LM (i.e., rationale learner ) to learn explicit denoising by leveraging these rationales as either in-context learning demonstrations or as supervised fine-tuning data (§ 2.3).
As detailed in Algorithm 1, during the entire process, InstructRAG does not require any additional supervisions beyond standard RAG methods. By default, we instantiate both and with the same off-the-shelf instruction-tuned model (i.e., meta-llama/Meta-Llama-3-8B-Instruct), making InstructRAG a fully self-synthesis method. We also experiment with different instantiations of and and conduct ablation study in both training-free and trainable settings (§ 3.3). For simplicity, we use placeholders to represent omitted instructions in the prompts presented in this section, while the full list of complete prompt templates is provided in Appendix C.
@@ 2.1 Problem Setting
We adopt the standard RAG setting where the LM has access to annotated datasets of downstream tasks (e.g., question-answering task ), and an external knowledge base with the off-the-shelf retriever for retrieval. Different from previous works [3, 105] which leverage additional supervisions from GPT-3 [6] or GPT-4 [1], we assume the model has strictly limited access to the above two information sources. Given a question , the retriever returns a set of potentially noisy documents from the external knowledge base. The model is then tasked to predict the correct answer to the given question based on and its own parametric knowledge, denoted as .
Our work focuses on investigating the noise robustness of LMs and developing efficient denoising techniques for RAG. Hence, we directly employ off-the-shelf retrievers instead of training our own, and prepend all retrieved documents to the question as input to the model, without any filtering or re-ranking. This setting is orthogonal to existing research efforts centered on optimizing the retriever or performing adaptive retrieval [3, 86, 102].
@@ 2.2 Rationale Generation via Instruction-Following
Recent studies [51, 63, 67] have made encouraging progress in aligning LMs with human preferences and intentions, enabling the synthesis of high-quality data that closely follows user instructions [98]. Inspired by these advances, we propose to leverage the LM’s strong instruction-following ability to generate explicit denoising responses (i.e., rationales) for RAG. As shown in Table 1, given a QA pair and a set of retrieved documents , we prompt an off-the-shelf LM (as the rationale generator) with denoising instructions to produce the corresponding rationale that distinguishes useful documents from noisy ones and explains how the contexts lead to the ground-truth answer . To ensure the synthetic rationales are aligned with the ground-truth answers, we use a simple substring match to assess their consistency. The consistency ratio on training samples with at least one relevant document containing the ground-truth answer is 98% on average across five benchmarks, supporting the reliability of synthetic rationales as a sanity check. This allows us to effectively augment the standard dataset with self-synthesized denoising rationales solely by instructing the LM, without any additional supervision.
@@ 2.3 Learning Denoising Rationales in RAG
InstructRAG-FT is a trainable instantiation of InstructRAG that learns denoising rationales via supervised fine-tuning (FT) with standard language modeling objective. As defined in Eq. (1), it maximizes the likelihood of rationale conditioned on question and retrieved documents .
@@ 3.1 Experimental Setting
RAG tasks and evaluation metrics. We extensively validate the effectiveness of InstructRAG on five knowledge-intensive benchmarks, including PopQA [61], TriviaQA [38], Natural Questions [47], ASQA [78], and 2WikiMultiHopQA [24]. We use Wikipedia corpus as the retrieval source, and test our method with both sparse and dense off-the-shelf retrievers, including BM25 [71], DPR [39], GTR [65] and Contriver [27]. The retrieval quality is measured by Recall@, indicating whether the retrieved documents contain the correct answer. Table 2 shows the detailed dataset statistics. Following standard evaluation settings [3], we adopt the official metric of correctness (str-em), citation precision (pre) and recall (rec) for ASQA [18], and use accuracy for the other tasks, which measures whether the ground-truth answers are included in the model generations [61, 73]. Additionally, we also adopt LLM-as-a-judge for further evaluation (§ 3.4), as the above standard metrics are subject to the limitations of pattern-matching, which cannot accurately handle semantic equivalence.
@@
Baselines. We compare our method with a wide range of RAG baselines under both training-free and trainable settings. Given that state-of-the-art LMs have incorporated a large amount of world-knowledge during the pre-training stage, we also report the performance of a non-retrieval baseline (namely, vanilla zero-shot prompting) for reference. Specifically, the training-free RAG baselines includes: (1) in-context retrieval-augmented language modeling (RALM) [70], a prompting method that extends the non-retrieval baseline by presenting the model with retrieved documents; (2) few-shot demonstration with instruction, an ICL method using ground-truth question-answer pairs sampled from the training set as demonstration exemplars.
@@
The trainable RAG baselines include: (1) vanilla supervised fine-tuning (SFT), a supervised method with the training objective of maximizing the data likelihood of ground-truth answer given potentially noisy input; (2) RetRobust [105], which fine-tunes the RAG model on a mixture of relevant and irrelevant contexts to make it robust to irrelevant contexts; (3) Self-RAG [3], a strong trainable baseline, focusing on adaptive retrieval controlled by special reflection tokens. Both RetRobust and Self-RAG were originally built on Llama-2 [82] with additional supervisions. For example, RetRobust augments the training data for multi-hop reasoning tasks (e.g., 2WikiMultiHopQA) by prompting GPT-3 to decompose the original query and generate intermediate subqueries, and Self-RAG requires GPT-4 to generate additional reflective tokens to augment training samples.
@@
For a fair comparison, we re-implement the two methods on Llama-2 and/or Llama-2 with augmented training data released by their authors, and report their performance as the higher one between the original scores and our reproduced results. As our method adopts instruction-tuned Llama-3 as the backbone model, we also train RetRobust and Self-RAG with Llama-3-Instruct and optimize their performance through extensive hyper-parameters search. More details on implementation, including training, inference, and prompt design are available in Appendix A and Appendix C.
@@ 3.2 Main Result
Baselines without retrieval. As shown in the first block, the basic instruction-tuned models (Llama-3-Instruct and Llama-3-Instruct) already achieve notable performance across all five benchmarks, with the 70B model exhibiting a surprisingly competitive performance of 80.6% on the TriviaQA. This observation suggests that the required knowledge for these tasks mostly falls within the LM’s parametric knowledge, probably due to what is known as data contamination (i.e., the presence of test data of downstream tasks in the pre-training data of LMs) [21, 30, 59].
@@
RAG without training. The second block shows the comparison among training-free RAG methods. In-context RALM and few-shot demonstration with instruction methods generally achieve higher performance than the non-retrieval baseline, highlighting the importance of retrieval for knowledge-intensive tasks. Encouragingly, our InstructRAG-ICL consistently outperforms all training-free baselines across various metrics, confirming the effectiveness of self-synthesized denoising rationales. Moreover, the boost from 8B to 70B model also demonstrates that InstructRAG-ICL scales effectively with larger backbone models, validating the generalizability of our method.
RAG with training. As present in the bottom block of Table 3, our InstructRAG-FT not only surpasses all non-retrieval and training-free baselines across all five benchmarks, but also significantly outperforms trainable RAG baselines on almost every metric. The only exception is in the ASQA task, where our method slightly underperforms Self-RAG in terms of citation (i.e., pre and rec). This is because our work primarily focuses on explicit denoising for RAG to improve the correctness of generations, which is measured by em. Despite not being explicitly optimized for citation metrics, our method still achieves competitive citation performance, significantly enhancing both generation accuracy and trustworthiness. Note that RetRobust achieves competitive performance on 2WikiMultiHopQA, which involves multi-hop reasoning. We attribute this to the additional training supervision provided by GPT-3, which enables the model to explicitly generate intermediate sub-queries and sub-answers. Another interesting finding is that Self-RAG consistently exhibits inferior performance compared to vanilla SFT, and even underperforms the training-free in-context RALM baseline across all benchmarks. We speculate the reason might be that these RAG tasks favor more domain-specific knowledge than general knowledge. However, it is challenging for Self-RAG to directly leverage in-domain features from existing training data as it requires GPT-4 to generate reflection tokens on these benchmarks, which is not available in our problem (§ 2.1).

### v3

# InstructRAG: Instructing Retrieval Augmented Generation via Self-Synthesized Rationales

@@ Abstract
Retrieval-augmented generation (RAG) has shown promising potential to enhance the accuracy and factuality of language models (LMs). However, imperfect retrievers or noisy corpora can introduce misleading or even erroneous information to the retrieved contents, posing a significant challenge to the generation quality. Existing RAG methods typically address this challenge by directly predicting final answers despite potentially noisy inputs, resulting in an implicit denoising process that is difficult to interpret and verify. On the other hand, the acquisition of explicit denoising supervision is often costly, involving significant human efforts. In this work, we propose InstructRAG, where LMs explicitly learn the denoising process through self-synthesized rationales — First, we instruct the LM to explain how the ground-truth answer is derived from retrieved documents. Then, these rationales can be used either as demonstrations for in-context learning of explicit denoising or as supervised fine-tuning data to train the model. Compared to standard RAG approaches, InstructRAG requires no additional supervision, allows for easier verification of the predicted answers, and effectively improves generation accuracy. Experiments show InstructRAG consistently outperforms existing RAG methods in both training-free and trainable scenarios, achieving a relative improvement of 8.3% over the best baseline method on average across five knowledge-intensive benchmarks. Extensive analysis indicates that InstructRAG scales well with increased numbers of retrieved documents and consistently exhibits robust denoising ability even in out-of-domain datasets, demonstrating strong generalizability.111Code is available at https://github.com/weizhepei/InstructRAG.
@@ 1 Introduction
While large language models (LMs) have demonstrated remarkable text generation abilities (Brown et al., 2020; Team et al., 2023; Touvron et al., 2023), they may occasionally produce factually incorrect contents (Dhuliawala et al., 2023; Huang et al., 2023a; Ji et al., 2023; Sun et al., 2023; Xu et al., 2024d; Zhang et al., 2023), particularly when the task at hand requires the most current information or out-of-domain knowledge not adequately represented in the pre-training corpus (Jiang et al., 2023b; Shuster et al., 2021; Yu et al., 2023; Zhao et al., 2023). This limitation significantly hinders the reliable deployment of LMs in high-stakes domains where factuality is crucial (Magesh et al., 2024; Singhal et al., 2023; Xiao et al., 2021; Xiong et al., 2024).
@@
In light of this, retrieval-augmented generation (RAG) (Asai et al., 2023b; Guu et al., 2020; Izacard et al., 2023; Khandelwal et al., 2019; Lewis et al., 2020) has been introduced to enhance the generation accuracy of LMs in knowledge-intensive tasks by leveraging the most up-to-date information and specialized knowledge from external sources (Kasai et al., 2024; Vu et al., 2023; Yang et al., 2024; Zhou et al., 2022). However, the retrieved contents are typically mixed with irrelevant or even erroneous information due to the absence of perfect retrieval solutions (Izacard et al., 2021; Karpukhin et al., 2020; Khattab et al., 2022; 2023; Shi et al., 2023; Su et al., 2024) and the presence of noisy data in the retrieval corpus (Izacard & Grave, 2021; Li et al., 2023; Yoran et al., 2024), posing a long-standing challenge to almost all RAG systems. Typically, vanilla RAG approaches address this issue implicitly by training LMs to directly predict correct answers despite noisy inputs. Such latent processes are not only difficult to interpret and verify but also vulnerable to higher noise ratios, especially when the number of retrieved documents is large (Chen et al., 2024; Cuconasu et al., 2024; Liu et al., 2024a; Wu et al., 2024). On the other hand, obtaining high-quality explicit denoising supervision often requires substantial human efforts, which is time-consuming and costly.
@@
In this work, we introduce a new RAG framework, InstructRAG, which enables the LM to explicitly denoise retrieved information and justify its predicted final answers by generating denoising responses (i.e., rationales), as illustrated in Figure 1. Compared to vanilla RAG approaches, InstructRAG does not require any additional supervision, while enjoying improved generation accuracy and trustworthiness. Specifically, our method consists of two steps. First, given a set of question-answer pairs and potentially noisy retrieved documents, we prompt an instruction-tuned LM to synthesize denoising rationales that analyze the documents and articulate how they lead to the ground-truth answers (§ 2.2). Then, these synthetic rationales can be utilized as in-context learning examples or as supervised fine-tuning data, allowing the LM to explicitly learn to denoise retrieved contents (§ 2.3). The effectiveness of InstructRAG can be attributed to the strong instruction-following ability of LMs (Jiang et al., 2024b; Ouyang et al., 2022; Wei et al., 2021), a significant feature that still remains underexplored in the context of RAG. We show that such self-synthesized rationales not only provide high-quality explicit denoising supervision for in-domain RAG tasks, but also facilitate superior out-of-domain generalization. This finding underscores how instruction-tuned LMs can synthesize generalizable supervision to overcome the inevitable noise in RAG.
@@ 2 Our Method: InstructRAG
In this section, we first introduce our problem setting (§ 2.1) and then present the proposed framework InstructRAG that enables LMs to explicitly denoise retrieved contents. As shown in Figure 2, our method consists of two steps. First, we prompt an instruction-tuned LM (i.e., rationale generator ) to synthesize rationales that provide denoising supervisions (§ 2.2). These rationales aim to explain how to derive the correct answer from potentially noisy retrieved documents for each training sample. Then, we guide the LM (i.e., rationale learner ) to learn explicit denoising by leveraging these rationales as either in-context learning demonstrations or as supervised fine-tuning data (§ 2.3). As detailed in Algorithm 1, during the entire process, InstructRAG does not require any additional supervisions beyond standard RAG methods. By default, we instantiate both and with the same off-the-shelf instruction-tuned model (i.e., meta-llama/Meta-Llama-3-8B-Instruct), making InstructRAG a fully self-synthesis method. We also experiment with different instantiations of and and conduct ablation study in both training-free and trainable settings (§ 3.3). For simplicity, we use placeholders to represent omitted instructions in the prompts presented in this section, while the full list of complete prompt templates is provided in Appendix D.
@@ 2.1 Problem Setting
We adopt the standard RAG setting where the LM has access to annotated datasets of downstream tasks (e.g., question-answering task ), and an external knowledge base with the off-the-shelf retriever for retrieval. Different from previous works (Asai et al., 2023b; Yoran et al., 2024) which leverage additional supervisions from GPT-3 (Brown et al., 2020) or GPT-4 (Achiam et al., 2023), we assume the model has strictly limited access to the above two information sources. Given a question , the retriever returns a set of potentially noisy documents from the external knowledge base. The model is then tasked to predict the correct answer to the given question based on and its own parametric knowledge, denoted as .
Our work focuses on investigating the noise robustness of LMs and developing efficient denoising techniques for RAG. Hence, we directly employ off-the-shelf retrievers instead of training our own, and prepend all retrieved documents to the question as input to the model, without any filtering or re-ranking. This setting is orthogonal to existing research efforts centered on optimizing the retriever or performing adaptive retrieval (Asai et al., 2023b; Wang et al., 2024a; Yang et al., 2024).
@@ 2.2 Rationale Generation via Instruction-Following
Recent studies (Leike et al., 2018; Meng et al., 2024; Ouyang et al., 2022) have made encouraging progress in aligning LMs with human preferences and intentions, enabling the synthesis of high-quality data that closely follows user instructions (Xu et al., 2024c). Inspired by these advances, we propose to leverage the LM’s strong instruction-following ability to generate explicit denoising responses (i.e., rationales) for RAG. As shown in Table 1, given a QA pair and a set of retrieved documents , we prompt an off-the-shelf LM (as the rationale generator) with denoising instructions to produce the corresponding rationale that distinguishes useful documents from noisy ones and explains how the contexts lead to the ground-truth answer . To ensure the synthetic rationales are aligned with the ground-truth answers, we use a simple substring match to assess their consistency. The consistency ratio on training samples with at least one relevant document containing the ground-truth answer is 98% on average across five benchmarks, supporting the reliability of synthetic rationales as a sanity check. This allows us to effectively augment the standard dataset with self-synthesized denoising rationales solely by instructing the LM, without any additional supervision.
@@ 2.3 Learning Denoising Rationales in RAG
InstructRAG-FT is a trainable instantiation of InstructRAG that learns denoising rationales via supervised fine-tuning (FT) with standard language modeling objective. As defined in Eq. 1, it maximizes the likelihood of rationale conditioned on question and retrieved documents .
@@ 3.1 Experimental Setting
RAG tasks and evaluation metrics. We extensively validate the effectiveness of InstructRAG on five knowledge-intensive benchmarks, including PopQA (Mallen et al., 2023), TriviaQA (Joshi et al., 2017), Natural Questions (Kwiatkowski et al., 2019), ASQA (Stelmakh et al., 2022), and 2WikiMultiHopQA (Ho et al., 2020). We use Wikipedia corpus as the retrieval source, and test our method with both sparse and dense off-the-shelf retrievers, including BM25 (Robertson & Walker, 1994), DPR (Karpukhin et al., 2020), GTR (Ni et al., 2022) and Contriver (Izacard et al., 2021). The retrieval quality is measured by Recall@, indicating whether the retrieved documents contain the correct answer. Table 2 shows the detailed dataset statistics. Following standard evaluation settings (Asai et al., 2023b), we adopt the official metric of correctness (str-em), citation precision (pre) and recall (rec) for ASQA (Gao et al., 2023a), and use accuracy for the other tasks, which measures whether the ground-truth answers are included in the model generations (Mallen et al., 2023; Schick et al., 2023). Additionally, we also adopt LLM-as-a-judge for further evaluation (§ 3.4), as the above standard metrics are subject to the limitations of pattern-matching, which cannot accurately handle semantic equivalence.
@@
Baselines. We compare our method with a wide range of RAG baselines under both training-free and trainable settings. Given that state-of-the-art LMs have incorporated a large amount of world-knowledge during the pre-training stage, we also report the performance of a non-retrieval baseline (namely, vanilla zero-shot prompting) for reference. Specifically, the training-free RAG baselines includes: (1) in-context retrieval-augmented language modeling (RALM) (Ram et al., 2023), a prompting method that extends the non-retrieval baseline by presenting the model with retrieved documents; (2) few-shot demonstration with instruction, an ICL method using ground-truth question-answer pairs sampled from the training set as demonstration exemplars.
@@
The trainable RAG baselines include: (1) vanilla supervised fine-tuning (SFT), a supervised method with the training objective of maximizing the data likelihood of ground-truth answer given potentially noisy input; (2) RetRobust (Yoran et al., 2024), which fine-tunes the RAG model on a mixture of relevant and irrelevant contexts to make it robust to irrelevant contexts; (3) Self-RAG (Asai et al., 2023b), a strong trainable baseline, focusing on adaptive retrieval controlled by special reflection tokens. Both RetRobust and Self-RAG were originally built on Llama-2 (Touvron et al., 2023) with additional supervisions. For example, RetRobust augments the training data for multi-hop reasoning tasks (e.g., 2WikiMultiHopQA) by prompting GPT-3 to decompose the original query and generate intermediate subqueries, and Self-RAG requires GPT-4 to generate additional reflective tokens to augment training samples.
@@
For a fair comparison, we re-implement the two methods on Llama-2 and/or Llama-2 with augmented training data released by their authors, and report their performance as the higher one between the original scores and our reproduced results. As our method adopts instruction-tuned Llama-3 as the backbone model, we also train RetRobust and Self-RAG with Llama-3-Instruct and optimize their performance through extensive hyper-parameters search. More details on implementation, including training, inference, and prompt design are available in Appendix B and Appendix D. We also present some case studies in Appendix C.
@@ 3.2 Main Result
Baselines without retrieval. As shown in the first block, the basic instruction-tuned models (Llama-3-Instruct and Llama-3-Instruct) already achieve notable performance across all five benchmarks, with the 70B model exhibiting a surprisingly competitive performance of 80.6% on the TriviaQA. This observation suggests that the required knowledge for these tasks mostly falls within the LM’s parametric knowledge, probably due to what is known as data contamination (i.e., the presence of test data of downstream tasks in the pre-training data of LMs) (Golchin & Surdeanu, 2023; Jacovi et al., 2023; Magar & Schwartz, 2022).
@@
RAG without training. The second block shows the comparison among training-free RAG methods. In-context RALM and few-shot demonstration with instruction methods generally achieve higher performance than the non-retrieval baseline, highlighting the importance of retrieval for knowledge-intensive tasks. Encouragingly, our InstructRAG-ICL consistently outperforms all training-free baselines across various metrics, confirming the effectiveness of self-synthesized denoising rationales. Moreover, the boost from 8B to 70B model indicates that InstructRAG-ICL scales effectively with larger backbone models, validating the generalizability of our method.
RAG with training. As present in the bottom block of Table 3, our InstructRAG-FT not only surpasses all non-retrieval and training-free baselines across all five benchmarks, but also significantly outperforms trainable RAG baselines on almost every metric. The only exception is in the ASQA task, where our method slightly underperforms Self-RAG in terms of citation (i.e., pre and rec). This is because our work primarily focuses on explicit denoising for RAG to improve the correctness of generations, which is measured by em. Despite not being explicitly optimized for citation metrics, our method still achieves competitive citation performance, significantly enhancing both generation accuracy and trustworthiness. Note that RetRobust achieves competitive performance on 2WikiMultiHopQA, which involves multi-hop reasoning. We attribute this to the additional training supervision provided by GPT-3, which enables the model to explicitly generate intermediate sub-queries and sub-answers. Another interesting finding is that Self-RAG consistently exhibits inferior performance compared to vanilla SFT, and even underperforms the training-free in-context RALM baseline across all benchmarks. We speculate the reason might be that these RAG tasks favor more domain-specific knowledge than general knowledge. However, it is challenging for Self-RAG to directly leverage in-domain features from existing training data as it requires GPT-4 to generate reflection tokens on these benchmarks, which is not available in our problem setting (§ 2.1).